# Always Keep Your Promises: A Model-Agnostic Attribution Algorithm for Neural Networks

### v3

@@ C.2 Promise Trees and Resolution
In this example, A and B are promise-generating operations, marked purple, and D, E, C are Arg Nodes, marked orange.
We also use "->" to indicate the node at which the traversal pointer is pointing at each step.
@@
We mark nodes as red when we have traversed them, blue when we have propagated relevance through them. It is important to note the difference between these two states due to the delaying of propagation through Promises. Node A is traversed first, and propagates Promise Branches $b_{A,1}$ and $b_{A,2}$ through to its out-neighbours.
@@ Appendix G Extra Results
VGG16 Simonyan and Zisserman (2015) Vision 90/90 9/9
ResNet-50 He et al. (2015) Vision 339/339 10/10
ViT-b-16 Dosovitskiy et al. (2021) Vision 779/779 16/16
EfficientNetv2-medium Tan and Le (2021) Vision 1,526/1,526 12/12
SigLIP-2-So400m-base14-384 Tschannen et al. (2025) Vision 2,178/2,178 19/19
Wav2Vec2-xls-r-300m Babu et al. (2021) Audio 2,021/2,022 18/19
whisper-small Radford et al. (2022) Audio 2,729/2,729 22/22
GPT-2 Radford et al. (2019) Language 885/885 19/19
RoBERTa-large Liu et al. (2019) Language 1,461/1,461 14/14
Llama3.2-1B Grattafiori et al. (2024) Language 1,787/1,787 24/24
Gemma-3-270m-it Gemma Team (2025) Language 2,482/2,482 24/24
Qwen3-0.6B Qwen Team (2025) Language 2,590/2,590 18/18
Flan-T5-large Chung et al. (2022) Language 5,713/5,713 24/24
DePlot Liu et al. (2022) Vision/Language 2,863/2,863 23/23
Mamba-130m Gu and Dao (2024) State Space 3,421/3,421 26/26

### v4

@@ C.2 Promise Trees and Resolution
In this example, A and B are promise-generating operations, marked purple, and D, E, C are Arg Nodes, marked orange. We also use "->" to indicate the node at which the traversal pointer is pointing at each step. We mark nodes as red when we have traversed them, blue when we have propagated relevance through them. It is important to note the difference between these two states due to the delaying of propagation through Promises. Node A is traversed first, and propagates Promise Branches $b_{A,1}$ and $b_{A,2}$ through to its out-neighbours.
@@ Appendix G Extra Results
VGG16 simonyan2015vgg Vision 90/90 9/9
ResNet-50 he2015resnet Vision 339/339 10/10
ViT-b-16 dosovitskiy2021vit Vision 779/779 16/16
EfficientNetv2-medium tan2021efficientnetv2 Vision 1,526/1,526 12/12
SigLIP-2-So400m-base14-384 tschannen2025siglip2 Vision 2,178/2,178 19/19
Wav2Vec2-xls-r-300m babu2021xlsr Audio 2,021/2,022 18/19
whisper-small radford2022whisper Audio 2,729/2,729 22/22
GPT-2 radford2019gpt2 Language 885/885 19/19
RoBERTa-large liu2019roberta Language 1,461/1,461 14/14
Llama3.2-1B dubey2024llama3 Language 1,787/1,787 24/24
Gemma-3-270m-it gemmateam2025gemma3 Language 2,482/2,482 24/24
Qwen3-0.6B yang2025qwen3 Language 2,590/2,590 18/18
Flan-T5-large chung2022flant5 Language 5,713/5,713 24/24
DePlot liu2022deplot Vision/Language 2,863/2,863 23/23
Mamba-130m gu2024mamba State Space 3,421/3,421 26/26
@@
Table A.6: Parameter randomization sanity check, as outlined in adebayo2020sanitycheckssaliencymaps, for DynamicLRP ViT-B-16 attributions on 100 CIFAR10 image classification examples. Layer randomization is done from output to input layer. For each sample, the Spearman correlation ($\rho$) was taken between the attribution produced by the trained model $M_{0}$ and each randomized model variant $M_{i}$, then the average over all 100 samples was taken to evaluate attribution deterioration throughout the randomization process.
Layers Randomized ($i$) Mean $\rho_{0,i}$
1 0.479
2 0.297
3 0.127
4 0.0239
5 0.0391
6 0.0310
7 0.0414
8 0.0214
9 0.0204
10 0.0260
11 0.00678
12 0.00785